OpenStudy (anonymous):

Use differentiation to find a power series representation for the following function. \[\frac{1}{(x+6)^2}\]

OpenStudy (blockcolder):

In any case, let's see what we can do. \[f'(x)=\frac{-2}{(x+6)^3}\] \[f''(x)=\frac{6}{(x+6)^4}\] \[f'''(x)=\frac{-24}{(x+6)^5}\] Can you see the pattern for the nth derivative? Also, where should the power series be centered?

OpenStudy (anonymous):

It should be centered at 0 and the nth derivative im not to sure on

OpenStudy (anonymous):

just taking a stab at the nth derivative.. is it \[\frac{(n+2)!}{(x+6)^{n+3}}\]

OpenStudy (blockcolder):

\[f^{(n)}(x)=\frac{(-1)^n (n+1)!}{(x+6)^{n+2}}\] Thus, \[c_n=\frac{f^{(n)}(0)}{n!}=\frac{(-1)^n n}{6^{n+2}}\] and its Taylor series is \[\sum_{n=0}^{\infty} \frac{(-1)^nn}{6^{n+2}}x\]

OpenStudy (anonymous):

ohh i started at n=0 oops

OpenStudy (blockcolder):

Alternatively, you can see that \[\frac{1}{(x+6)^2}={dy \over dx}\left (\frac{-1}{x+6}\right)\]. \[\frac{-1}{x+6}=\frac{1}{6-x}\]. Do you know how to get the power series of 1/(6-x)?

OpenStudy (blockcolder):

Err, I meant 1/(-6-x), not 1/(6-x). :))

OpenStudy (anonymous):

not quite

OpenStudy (blockcolder):

Anyway, the power series of -1/(x+6) is: \[\begin{align} \frac{-1}{x+6} & =\frac{-1}{6} \cdot \frac{1}{\frac{x}{6}+1} \\ & = {-1 \over 6} \sum_{n=0}^{\infty} \left ( {-x \over 6} \right )^{n} \\ & = {-1 \over 6} \sum_{n=0}^{\infty} {(-1)^n x^n \over 6^n} \end{align}\]

OpenStudy (blockcolder):

That is equal to \[\sum_{n=0}^{\infty} {(-1)^{n+1} x^n \over6^{n+1}}\] then differentiate this with respect to x.

OpenStudy (blockcolder):

Sorry for the late reply and any errors you see. I'm very tired.

OpenStudy (anonymous):

Its ok, thank you

OpenStudy (blockcolder):

And as I typed my previous reply, I saw an error in my first answer. n should be n+1 and x should be x^n. Anyway, the second solution should produce the same answer as the first.